# An Efficient Local Search Approach for Polarized Community Discovery in Signed Networks

### v2

@@ 4 Experiments
In this section, we present our experimental evaluation. Additional results are provided in Appendix G. We use eight publicly available real-world datasets commonly adopted in prior work on PCD DBLP:conf/nips/TzengOG20 , with dataset details included in Appendix G.1. Notably, no real-world datasets with ground-truth solutions for PCD (i.e., where neutral objects are allowed) currently exist. Consequently, prior work has relied on polarity (Eq. 2) as a proxy for evaluating solution quality DBLP:conf/cikm/BonchiGGOR19 ; DBLP:conf/nips/TzengOG20 . For consistency, we also report polarity scores for these datasets. Additionally, following previous work DBLP:conf/nips/TzengOG20 , we include experiments on synthetic datasets where ground-truth solutions are available. Throughout this section, we fix $\alpha$ and $\beta$ (as specified below). In Appendix C, we provide a detailed discussion on the impact of these parameters (and we investigate it experimentally in Appendix G.7). We compare our local search algorithm for PCD, named LSPCD, to several baseline methods, which we introduce below.
@@
Metrics. (i) Following prior work, we use polarity to evaluate the quality of different methods DBLP:conf/cikm/BonchiGGOR19 ; DBLP:conf/nips/TzengOG20 , defined as in Eq. 2 with $\alpha=1/(k-1)$. (ii) For datasets with available ground-truth, we measure the recovery-rate of ground-truth clusters using the F1-score, which is the precision and recall averaged over all clusters (as in (DBLP:conf/cikm/BonchiGGOR19, ; DBLP:conf/nips/TzengOG20, )). (iii) To evaluate the balance of a clustering solution $S_{[k]}$, we use the imbalance factor from (DBLP:journals/kais/PirizadehFK23, ). Let $p_{i}=|S_{i}|/\sum_{m\in[k]}|S_{m}|$ be the proportion of objects in cluster $S_{i}$. The imbalance factor is defined as $\frac{1}{1-\xi}\log_{2}(\sum_{i=1}^{k}p_{i}^{\xi})/\log_{2}(k)\in[0,1]$, where 1 indicates perfect balance and 0 indicates maximal imbalance (i.e., all objects in one cluster). For $\xi=1$, the numerator reduces to Shannon entropy; we use $\xi=3$ to penalize highly imbalanced solutions more strongly. Appendix G.8 presents a detailed summary of the solutions found by each method, including the number of non-neutral objects, the number of non-empty clusters, runtime, and more.
Synthetic datasets. In our first experiment, we evaluate how well the methods recover ground-truth clusters using synthetic networks. Following DBLP:conf/cikm/BonchiGGOR19 ; DBLP:conf/nips/TzengOG20 , we employ the modified signed stochastic block model (m-SSBM), which was specifically designed to generate synthetic graphs with planted ground-truth communities for PCD. The m-SSBM model is parameterized by four variables: (i) $n$, the total number of nodes; (ii) $k$, the number of non-neutral clusters; (iii) $\ell$, the size of each non-neutral cluster; and (iv) $\eta\in[0,1]$, which controls the edge probabilities. Smaller values of $\eta$ correspond to denser non-neutral clusters and lower levels of noise (see Appendix G.3 for detailed description). Here, we assume balanced ground-truth clusters of size $\ell$. In Appendix G.4, we show that our method remains robust when increasing cluster imbalance on synthetic data. In Fig. 1, we present the F1-score and polarity of various methods on different synthetic graphs generated using the m-SSBM model, across different noise levels $\eta$. For clarity, we include the best performing baselines (KOCG peforms poorly here). Each setting is repeated 10 times, and we report the average. We fix $\beta=0.4$ for LSPCD (it is robust to the choice of $\beta$) and $\gamma=1$ for N2PC. We see that the recovery rate of all methods decline as $\eta$ increases, since the sparsity and noise level of the graph increases. For $k=4,6,10$, we observe that our method significantly outperforms baseline methods, being the only method capable of recovering the ground-truth solutions for $\eta>0.2$. For $k=2$, we observe that our method, SCG-MA, and N2PC perform the best. However, for $k=2$ (which N2PC is limited to), the problem is significantly more simple. Finally, we see that the ground-truth solution correlates with large polarity, justifying the use of polarity to measure solution quality for the real-world data.
@@
Real-world datasets. Table 1 present results for different methods and datasets with $k=2,4,6$ (we use the same values of $k$ as SCG DBLP:conf/nips/TzengOG20 ). $|E|$ denotes the number of edges with non-zero edge weight. Only seven of the eight datasets are shown (due to space limit); see Appendix G.8 for complete results. The spectral clustering methods, BNC and SPONGE, exceeded memory limits on large datasets (caused by $k$-means), indicated by dashes. We report the mean over five runs, with standard deviations included in Appendix G.8. For our method, we select the $\beta$ value that maximizes polarity, testing 10 values per dataset, while we fix $\alpha=1/(k-1)$ for all methods. For each method and dataset, we report the polarity (POL) and the imbalance factor (IF). Results show that our method is highly competitive—often the best—in polarity across all datasets. In particular, our method consistently finds solutions with large polarity, while maintaining a good cluster size balance (large imbalance factor). Additionally, our method does not impose strict balance constraints, which is beneficial since real-world clusterings are rarely perfectly balanced. Instead, it identifies high-polarity solutions with reasonable balance, making it more practical for real-world applications. Notably, in cases where baseline methods attain higher polarity, it is usually at the cost of a very low imbalance factor (which often implies one or more empty clusters). Moreover, baseline methods with an imbalance factor near 1 generally exhibit very low polarity. This observation highlights the inherent trade-off between polarity and cluster balance—which our approach balances very well. Furthermore, in Table 2, we show results comparing our method with N2PC for $k=2$. We observe that increasing $\gamma$ results in more balanced clusters, which consistently leads to lower polarity (consistent with Table 1). When N2PC optimizes standard polarity ($\gamma=1$), the polarity is usually highest, but the imbalance factor is consistently very low (often near zero). This suggests that optimizing polarity alone (as SCG does) is not ideal; encouraging the algorithm to produce more balanced solutions (even at the cost of reduced polarity) generally yields solutions that align better with user expectations in practice (see Appendix E for a detailed discussion on this). This is a well-known observation in previous work on clustering of signed and unsigned graphs (beyond PCD, without neutral objects) DBLP:conf/cikm/ChiangWD12 . While N2PC is competitive with our method, it is limited to $k=2$, requires tuning $\gamma$, and is significantly more complex (it requires training a GNN, leading to higher runtime, see Appendix G.8).
Table 1: Polarity (POL) and imbalance factor (IF) for different methods and real-world datasets.
@@
Table 2: Polarity (POL) and imbalance factor (IF) for LSPCD and N2PC ($k=2$).
BTC WikiV REF SD WikiC EP WikiP
$k$ POL IF POL IF POL IF POL IF POL IF POL IF POL IF
@@ 5 Conclusion
We proposed a novel formulation of the polarized community discovery (PCD) problem that emphasizes (reasonably) balanced communities in terms of size—addressing a key limitation of prior work, which typically optimizes polarity (Eq. 2) and often produces highly imbalanced clusterings. To tackle this, we developed the first efficient and scalable local search method for PCD and established a connection to block-coordinate Frank-Wolfe (FW) optimization. While the standard FW algorithm is known to achieve a convergence rate of $O(1/\sqrt{t})$ for general non-concave objectives DBLP:journals/corr/Lacoste-Julien16 ; DBLP:conf/allerton/ReddiSPS16 , we showed that—due to the specific structure of our objective in Eq. 5—our method achieves a significantly faster linear convergence rate of $O(1/t)$, despite the function being both non-concave and non-multilinear. Extensive experiments demonstrated that our method (LSPCD) consistently produces high-quality clusterings with reasonable cluster size balance, better aligning with practical expectations. Furthermore, we observed that the strong performance of local search algorithms in correlation clustering carried over to the PCD setting as well. Overall, our approach offers a compelling alternative in the PCD literature, both in terms of performance and simplicity (see Alg. 2). Alternative methods in the literature are significantly more complex.
@@ Appendix A Comparing CC and PCD
The following proposition presents alternative objectives equivalent to maximizing Eq. 1 (i.e., solving the $k$-CC problem, see Problem 1). While this is known in the CC literature [9], we include a complete summary here to better motivate our problem formulation of PCD.
@@
Problem 1 is equivalent to finding a clustering $S_{[k]}$ that maximizes any one of the four objectives in Eqs. 9-12 (i.e., they share all local maxima).
@@
$-N^{-}_{\text{intra}}-N^{+}_{\text{inter}}$ (10)
@@
A clustering $S_{[k]}$ with neutral objects $S_{0}=V\setminus\bigcup_{m\in[k]}S_{m}$ that maximizes one of the objectives in Eq. 1 or Eqs. 9-12 is not guaranteed to maximize any of the other objectives [^1] Unless $k=2$, in which case Eq. 9 and Eq. 1 are equivalent as established in [6]..
@@ Appendix B Proofs
Defining $A^{\prime}_{i,j}=(1+\alpha)(A_{i,j}-\beta)$, we observe that since $c_{\text{sim}}$ is a constant across clustering solutions, the problem reduces to finding a partition of $V^{\prime}$ that maximizes $\sum_{m\in[k]}\sum_{i,j\in S_{m}}A^{\prime}_{i,j}$. This is equivalent to the max correlation objective (Eq. 11) applied to the transformed adjacency matrix $A^{\prime}$. By Prop. 2, this objective is equivalent to the $k$-CC problem (Problem 1). Thus, solving the $k$-PCD problem requires solving the $k$-CC problem on the instance $G^{\prime}=(V^{\prime},E^{\prime})$, meaning $k$-PCD is at least as hard as $k$-CC. Since correlation clustering is NP-hard [5, 20], we conclude that $k$-PCD is also NP-hard. ∎
@@
The second equality (line 3) holds because the number of pairs of objects inside cluster $m$ is $|S_{m}|^{2}$. A similar regularization is established in [10] for the minimum cut objective, where it is shown that optimizing this minimum cut objective regularized with $-\beta\sum_{m\in[k]}|S_{m}|^{2}$ is equivalent to optimizing the max correlation objective (Eq. 11) with similarities shifted by $\beta$. However, their result specifically considers the full network partitioning of unsigned networks, where the initial pairwise similarities are assumed non-negative. Moreover, they use a different regularization in practice: they shift the pairwise similarities so that the sum of the rows and columns of the similarity matrix becomes zero. ∎
@@
$=-\beta+2\sum_{j\in S_{m}\setminus\{i\}}A_{i,j}-2|S_{m}|\beta+2\beta\bm{1}_{[i\in S_{m}]}-2\alpha\sum_{p\in[k]\setminus\{m\}}\sum_{j\in S_{p}\setminus\{i\}}A_{i,j}.$
In the last equality we use $-2\sum_{j\in S_{m}\setminus\{i\}}\beta=-2|S_{m}|\beta+2\beta\bm{1}_{[i\in S_{m}]}$. We note that computing the final expression in Eq. 38 is $O(k^{2}n)$, due to the last term. However, by noting that $\sum_{p\in[k]\setminus\{m\}}\sum_{j\in S_{p}\setminus\{i\}}A_{i,j}=\sum_{j\notin S_{0}}A_{i,j}-\sum_{j\in S_{m}\setminus\{i\}}A_{i,j}$ we can derive the following.
$c_{im}$ $=-\beta+2\sum_{j\in S_{m}\setminus\{i\}}A_{i,j}-2|S_{m}|\beta+2\beta\bm{1}_{[i\in S_{m}]}-2\alpha\sum_{p\in[k]\setminus\{m\}}\sum_{j\in S_{p}\setminus\{i\}}A_{i,j}$ (39)
$=-\beta+2\sum_{j\in S_{m}\setminus\{i\}}A_{i,j}-2|S_{m}|\beta+2\beta\bm{1}_{[i\in S_{m}]}-2\alpha\big{(}\sum_{j\notin S_{0}}A_{i,j}-\sum_{j\in S_{m}\setminus\{i\}}A_{i,j}\big{)}$
$=-\beta+2\sum_{j\in S_{m}\setminus\{i\}}(A_{i,j}+\alpha A_{i,j})-2|S_{m}|\beta+2\beta\bm{1}_{[i\in S_{m}]}-2\alpha\sum_{j\notin S_{0}}A_{i,j}$
$=-\beta+2(1+\alpha)\sum_{j\in S_{m}\setminus\{i\}}A_{i,j}-2|S_{m}|\beta+2\beta\bm{1}_{[i\in S_{m}]}-2\alpha\sum_{j\notin S_{0}}A_{i,j}$
Now, recall the defintions $M_{i,m}\coloneqq 2\sum_{j\in S_{m}}A_{i,j}$, $\phi_{i}\coloneqq\sum_{m\in[k]}M_{i,m}$ and $\beta_{im}=2\beta|S_{m}|-2\beta\bm{1}_{[i\in S_{m}]}$. Then, we note that $2\sum_{j\notin S_{0}}A_{i,j}=\sum_{m\in[k]}M_{i,m}$ (sum of all similarities from object $i$ to all non-neutral objects). Then, we obtain the final expression, i.e.,
$c_{im}=-\beta+(1+\alpha)M_{i,m}-\beta_{im}-\alpha\phi_{i},\quad\text{for }m\in[k],$ (40)
The expression in Eq. 40 can be computed in $O(kn)$ since it just amounts to computing $M_{i,m}$ for all $m\in[k]$ (i.e., the total similarity from $i$ to all non-neutral cluters $m\in[k]$).
@@
We begin by defining the following quantities, which are constants w.r.t. different clustering solutions for the $k$-CC problem.

### v3

@@ 4 Experiments
In this section, we present our experimental evaluation. Additional results are provided in Appendix G. We use eight publicly available real-world datasets commonly adopted in prior work on PCD DBLP:conf/nips/TzengOG20 , with dataset details included in Appendix G.1. Notably, no real-world datasets with ground-truth solutions for PCD (i.e., where neutral objects are allowed) currently exist. Consequently, prior work has relied on polarity (Eq. 2) as a proxy for evaluating solution quality DBLP:conf/cikm/BonchiGGOR19 ; DBLP:conf/nips/TzengOG20 . For consistency, we also report polarity scores for these datasets. Additionally, following previous work DBLP:conf/nips/TzengOG20 , we include experiments on synthetic datasets where ground-truth solutions are available. Throughout this section, we fix $\alpha$ and $\beta$ (as specified below). In Appendix C, we provide a detailed discussion on the impact of these parameters (and we investigate it experimentally in Appendix G.8). We compare our local search algorithm for PCD, named LSPCD (see Alg. 3), to several baseline methods, which we introduce below. The complete source code for all experiments is publicly available [^1] https://github.com/Linusaronsson/NeurIPS2025-LSPCD.
@@
Metrics. (i) Following prior work, we use polarity to evaluate the quality of different methods DBLP:conf/cikm/BonchiGGOR19 ; DBLP:conf/nips/TzengOG20 , defined as in Eq. 2 with $\alpha=1/(k-1)$. (ii) For datasets with available ground-truth, we measure the recovery-rate of ground-truth clusters using the F1-score, which is the precision and recall averaged over all clusters (as in (DBLP:conf/cikm/BonchiGGOR19, ; DBLP:conf/nips/TzengOG20, )). (iii) To evaluate the balance of a clustering solution $S_{[k]}$, we use the imbalance factor from (DBLP:journals/kais/PirizadehFK23, ). Let $p_{i}=|S_{i}|/\sum_{m\in[k]}|S_{m}|$ be the proportion of objects in cluster $S_{i}$. The imbalance factor (IF) is defined as
$\texttt{IF}(p_{1},\dots,p_{k})=\frac{1}{1-\xi}\log_{2}(\sum_{i=1}^{k}p_{i}^{\xi})/\log_{2}(k)\in[0,1],$ (10)
where 1 indicates perfect balance and 0 indicates maximal imbalance (i.e., all objects in one cluster). For $\xi=1$, the numerator reduces to Shannon entropy; we use $\xi=3$ to penalize highly imbalanced solutions more strongly. The conclusions of our results are robust to changes in $\xi$ around our chosen value. Results for other values of $\xi$ are provided in Appendix G.6. In addition, Appendix G.9 presents a detailed summary of the solutions found by each method, including the number of non-neutral objects, the number of non-empty clusters, runtime, and more.
@@
Synthetic datasets. In our first experiment, we evaluate how well the methods recover ground-truth clusters using synthetic networks. Following DBLP:conf/cikm/BonchiGGOR19 ; DBLP:conf/nips/TzengOG20 , we employ the modified signed stochastic block model (m-SSBM), which was specifically designed to generate synthetic graphs with planted ground-truth communities for PCD. The m-SSBM model is parameterized by four variables: (i) $n$, the total number of nodes; (ii) $k$, the number of non-neutral clusters; (iii) $\ell$, the size of each non-neutral cluster; and (iv) $\eta\in[0,1]$, which controls the edge probabilities. Smaller values of $\eta$ correspond to denser non-neutral clusters and lower levels of noise (see Appendix G.3 for detailed description). Here, we assume balanced ground-truth clusters of size $\ell$. In Appendix G.4, we show that our method remains robust when increasing cluster imbalance on synthetic data.
In Figure 2, we present the F1-score and polarity of various methods on different synthetic graphs generated using the m-SSBM model, across different noise levels $\eta$. For clarity, we include the best performing baselines (KOCG peforms poorly here). Each setting is repeated 10 times, and we report the average. We fix $\beta=0.4$ for LSPCD (it is robust to the choice of $\beta$) and $\gamma=1$ for N2PC. We see that the recovery rate of all methods decreases as $\eta$ increases, since the sparsity and noise level of the graph increases. For $k=4,6,10$, we observe that our method significantly outperforms baseline methods, being the only method capable of recovering the ground-truth solutions for $\eta>0.2$. For $k=2$, we observe that our method, SCG-MA, and N2PC perform the best. However, for $k=2$ (which N2PC is limited to), the problem is significantly more simple. Finally, we see that the ground-truth solution correlates with large polarity, justifying the use of polarity to measure solution quality for the real-world data.
Table 1: Polarity (POL) and imbalance factor (IF) for different methods and real-world datasets. $|E|$ denotes the number of edges with non-zero edge weight.
@@
Real-world datasets. Table 1 present results for different methods and datasets with $k=2,4,6$ (we use the same values of $k$ as SCG DBLP:conf/nips/TzengOG20 ). $|E|$ denotes the number of edges with non-zero edge weight. Only seven of the eight datasets are shown (due to space limit); see Appendix G.9 for complete results. The spectral clustering methods, BNC and SPONGE, exceeded memory limits on large datasets (caused by $k$-means), indicated by dashes. We report the mean over five runs, with standard deviations included in Appendix G.9. For our method, we select the $\beta$ value that maximizes polarity, testing 10 values per dataset, while we fix $\alpha=1/(k-1)$ for all methods. For each method and dataset, we report the polarity (POL) and the imbalance factor (IF).
The results show that our method is highly competitive, often the best, in polarity across all datasets. In particular, our method consistently finds solutions with large polarity, while maintaining a good cluster size balance (large imbalance factor). Additionally, our method does not impose strict balance constraints, which is beneficial since real-world clusterings are rarely perfectly balanced. Instead, it identifies high-polarity solutions with reasonable balance, making it more practical for real-world applications. Notably, in cases where baseline methods attain higher polarity, it is usually at the cost of a very low imbalance factor (which often implies one or more empty clusters). Moreover, baseline methods with an imbalance factor near 1 generally exhibit very low polarity. This observation highlights the inherent trade-off between polarity and cluster balance, which our approach balances very well.
Results for N2PC are only included for $k=2$, as it does not support $k>2$. We observe that increasing $\gamma$ results in more balanced clusters, which consistently leads to lower polarity. When N2PC optimizes standard polarity ($\gamma=1$), the polarity is usually highest, but the imbalance factor is consistently very low (often near zero). This suggests that optimizing polarity alone (as SCG does) is not ideal; encouraging the algorithm to produce more balanced solutions (even at the cost of reduced polarity) generally yields solutions that align better with user expectations in practice (see Appendix E for a detailed discussion on this). This is a well-known observation in previous work on clustering of signed and unsigned graphs (beyond PCD, without neutral objects) DBLP:conf/cikm/ChiangWD12 . While N2PC is competitive with our method, it is limited to $k=2$, requires tuning $\gamma$, and is significantly more complex (it requires training a graph neural network, leading to higher runtime, see Appendix G.9).
@@ 5 Conclusion
We proposed a novel formulation of the polarized community discovery (PCD) problem that emphasizes (reasonably) balanced communities in terms of size, addressing a key limitation of prior work, which typically optimizes polarity (Eq. 2) and often produces highly imbalanced clusterings. To tackle this, we developed the first efficient and scalable local search method for PCD and established a connection to block-coordinate Frank-Wolfe (FW) optimization. While the standard FW algorithm is known to achieve a convergence rate of $O(1/\sqrt{t})$ for general non-concave objectives DBLP:journals/corr/Lacoste-Julien16 ; DBLP:conf/allerton/ReddiSPS16 , we showed that, due to the specific structure of our objective in Eq. 6, our method achieves a significantly faster linear convergence rate of $O(1/t)$, despite the function being both non-concave and non-multilinear. Extensive experiments demonstrated that our method (LSPCD) consistently produces high-quality clusterings with reasonable cluster size balance, better aligning with practical expectations. Furthermore, we observed that the strong performance of local search algorithms in correlation clustering carried over to the PCD setting as well. Overall, our approach offers a compelling alternative in the PCD literature, both in terms of performance and simplicity (see Alg. 2). Alternative methods in the literature are significantly more complex.
@@ Appendix A Comparing CC and PCD
The following proposition presents alternative objectives equivalent to maximizing Eq. 1 (i.e., solving the $k$-$C$ problem, see Problem 1). While this is known in the CC literature [9], we include a complete summary here to better motivate our problem formulation of PCD.
@@
Problem 1 is equivalent to finding a clustering $S_{[k]}$ that maximizes any one of the four objectives in Eqs. 11-14 (i.e., they share all local maxima).
@@
$N^{-}_{\text{intra}}-N^{+}_{\text{inter}}$ (12)
@@
A clustering $S_{[k]}$ with neutral objects $S_{0}=V\setminus\bigcup_{m\in[k]}S_{m}$ that maximizes one of the objectives in Eq. 1 or Eqs. 11-14 is not guaranteed to maximize any of the other objectives [^2] Unless $k=2$, in which case Eq. 11 and Eq. 1 are equivalent as established in [6]..
@@ Appendix B Proofs
Defining $A^{\prime}_{i,j}=(1+\alpha)(A_{i,j}-\beta)$, we observe that since $c_{\text{sim}}$ is a constant across clustering solutions, the problem reduces to finding a partition of $V^{\prime}$ that maximizes $\sum_{m\in[k]}\sum_{i,j\in S_{m}}A^{\prime}_{i,j}$. This is equivalent to the max correlation objective (Eq. 13) applied to the transformed adjacency matrix $A^{\prime}$. By Prop. 2, this objective is equivalent to the $k$-$C$ problem (Problem 1). Thus, solving the $k$-PCD problem requires solving the $k$-$C$ problem on the instance $G^{\prime}=(V^{\prime},E^{\prime})$, meaning $k$-PCD is at least as hard as $k$-$C$. Since correlation clustering is NP-hard [5, 20], we conclude that $k$-PCD is also NP-hard. ∎
@@
The second equality (line 3) holds because the number of pairs of objects inside cluster $m$ is $|S_{m}|^{2}$. A similar regularization is established in [10] for the minimum cut objective, where it is shown that optimizing this minimum cut objective regularized with $\beta\sum_{m\in[k]}|S_{m}|^{2}$ is equivalent to optimizing the max correlation objective (Eq. 13) with similarities shifted by $\beta$. However, their result specifically considers the full network partitioning of unsigned networks, where the initial pairwise similarities are assumed non-negative. Moreover, they use a different regularization in practice: they shift the pairwise similarities so that the sum of the rows and columns of the similarity matrix becomes zero. ∎
@@
$=-\beta+2\sum_{j\in S_{m}\setminus\{i\}}A_{i,j}-2|S_{m}|\beta+2\beta\bm{1}[i\in S_{m}]-2\alpha\sum_{p\in[k]\setminus\{m\}}\sum_{j\in S_{p}\setminus\{i\}}A_{i,j}.$
In the last equality we use $2\sum_{j\in S_{m}\setminus\{i\}}\beta=-2|S_{m}|\beta+2\beta\bm{1}[i\in S_{m}]$. We note that computing the final expression in Eq. 40 is $O(k^{2}n)$, due to the last term. However, by noting that $\sum_{p\in[k]\setminus\{m\}}\sum_{j\in S_{p}\setminus\{i\}}A_{i,j}=\sum_{j\notin S_{0}}A_{i,j}-\sum_{j\in S_{m}\setminus\{i\}}A_{i,j}$ we can derive the following.
$c_{im}$ $=-\beta+2\sum_{j\in S_{m}\setminus\{i\}}A_{i,j}-2|S_{m}|\beta+2\beta\bm{1}[i\in S_{m}]-2\alpha\sum_{p\in[k]\setminus\{m\}}\sum_{j\in S_{p}\setminus\{i\}}A_{i,j}$ (41)
$=-\beta+2\sum_{j\in S_{m}\setminus\{i\}}A_{i,j}-2|S_{m}|\beta+2\beta\bm{1}[i\in S_{m}]-2\alpha\big(\sum_{j\notin S_{0}}A_{i,j}-\sum_{j\in S_{m}\setminus\{i\}}A_{i,j}\big)$
$=-\beta+2\sum_{j\in S_{m}\setminus\{i\}}(A_{i,j}+\alpha A_{i,j})-2|S_{m}|\beta+2\beta\bm{1}[i\in S_{m}]-2\alpha\sum_{j\notin S_{0}}A_{i,j}$
$=-\beta+2(1+\alpha)\sum_{j\in S_{m}\setminus\{i\}}A_{i,j}-2|S_{m}|\beta+2\beta\bm{1}[i\in S_{m}]-2\alpha\sum_{j\notin S_{0}}A_{i,j}$
Then, we note that $2\sum_{j\notin S_{0}}A_{i,j}=2\sum_{p\in[k]}\sum_{j\in S_{p}}A_{i,j}=\sum_{p\in[k]}M_{i,p}$ (sum of all similarities from object $i$ to all non-neutral objects) and that $A_{i,i}=0$ (by assumption). This proves the statement of the theorem. The expression in Eq. 9 can be computed in $O(kn)$ since $\sum_{p\in[k]}\sum_{j\in S_{p}}A_{i,j}$ is a constant w.r.t. different clusters $m\in[k]$, and can therefore be precomputed (see Alg. 3). It may appear reasonable to remove the term $\sum_{p\in[k]}\sum_{j\in S_{p}}A_{i,j}$, since it is constant with respect to the cluster $m\in[k]$. However, this is invalid because neutral objects are allowed. Specifically, if $c_{im}<0$ for all $m\in[k]$, assigning object $i$ to the neutral set is optimal, as $c_{i0}=0$. The term $\sum_{p\in[k]}\sum_{j\in S_{p}}A_{i,j}$ must therefore be retained, as it affects whether an object is assigned to the neutral set. ∎
@@
We begin by defining the following quantities, which are constants w.r.t. different clustering solutions for the $k$-$C$ problem.